There is a pole in a lake. 1/2 of the pole is in the ground, another 1/3 of it is covered with water, and 10ft is out of water. What is the toal length of the pole in ft?
Answer by stanbon(75887)

There is a pole in a lake. 1/2 of the pole is in the ground, another 1/3 of it is covered with water, and 10ft is out of water. What is the toal length of the pole in ft?
----------------
Let the length of the pole be "x" ft.
(1/2)x + (1/3)x = x-10
Multiply thru by 6 to get:
3x + 2x = 6x-60
-x = -60
x = 60 ft.
=================
Cheers,
Stan H.